# CAS-ViT: Convolutional Additive Self-attention Vision Transformers for Efficient Mobile Applications

### v1

@@ Abstract
Vision Transformers (ViTs) mark a revolutionary advance in neural networks with their token mixer’s powerful global context capability. However, the pairwise token affinity and complex matrix operations limit its deployment on resource-constrained scenarios and real-time applications, such as mobile devices, although considerable efforts have been made in previous works. In this paper, we introduce CAS-ViT: Convolutional Additive Self-attention Vision Transformers, to achieve a balance between efficiency and performance in mobile applications. Firstly, we argue that the capability of token mixers to obtain global contextual information hinges on multiple information interactions, such as spatial and channel domains. Subsequently, we construct a novel additive similarity function following this paradigm and present an efficient implementation named Convolutional Additive Token Mixer (CATM). This simplification leads to a significant reduction in computational overhead. We evaluate CAS-ViT across a variety of vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation. Our experiments, conducted on GPUs, ONNX, and iPhones, demonstrate that CAS-ViT achieves a competitive performance when compared to other state-of-the-art backbones, establishing it as a viable option for efficient mobile vision applications. Our code and model are available at: https://github.com/Tianfang-Zhang/CAS-ViT
Introduction
@@
In recent years, the emergence of Vision Transformers (ViTs) marks a revolutionary shift in neural network architecture (Han et al. 2021; Huang et al. 2023; Zhang et al. 2023). Compared to Convolutional Neural Networks (CNNs), which make a splash with lower computational complexity and higher inference efficiency (Zhang et al. 2022; Zhang, Li, and Peng 2023), ViTs employ a novel architecture known as token mixer (Yu et al. 2022). This module facilitates enhanced global modeling and representational capabilities of ViTs through its ability to capture long-range dependencies.
The basic module of ViT consists of token mixer, MLP, and corresponding skip connections, where token mixer is widely implemented as Multi-head Self-Attention (MSA). MSA operates on the entire input sequence, compensating for the limitations of CNNs with restricted receptive fields, and providing unique advantages in terms of model scalability and adaptability (Li et al. 2023; Wu et al. 2024). Despite its reputation, the reality that MSA is of quadratic complexity with respect to input image size (Wu et al. 2021) makes ViT models more resource-intensive, rending it unsuitable for real-time applications and unfavorable for widespread deployment on resource-constrained devices, e.g. mobile applications (Mehta and Rastegari 2021; Liu et al. 2023). Therefore, developing token mixers that combine both efficient deployment and high performance becomes an urgent problem on mobile devices.
@@
Efforts have been made to improve token mixers, including MSA refinements and Heterogeneous-MSA (H-MSA). Refinements focus on following matrix multiplication of Query and Key to enhance the ability to capture long-range dependencies while reducing algorithmic complexity. Specific techniques include feature shift (Liu et al. 2021), carrier token (Hatamizadeh et al. 2023), sparse attention (Wang et al. 2022), linear attention (Han et al. 2023), etc. As an alternative paradigm, H-MSA intends to break through the limitations of point-by-point similarity between Q and K to explore more flexible network designs (Yu et al. 2022). The recently proposed pool token mixer (Yu et al. 2022) and context vector (Sandler et al. 2018; Shaker et al. 2023) further accelerate the inference efficiency.
@@
Although considerable progress has been made in related works, ViT models still suffer from the following constraints: 1. High complexity associated with matrix operations such as Softmax in token mixers. 2. The difficulty of simultaneously achieving accuracy, efficiency and easy-to-deployment on mobile devices or real-time applications.
In order to address the encountered problems, we propose a family of lightweight networks named as Convolutional Additive Self-attention (CAS)-ViT to achieve a balance between computation and efficiency as shown in Fig.1. Firstly, we argue that the capability of token mixers to obtain global contextual information hinges on multiple information interactions, such as spatial and channel domains. Meanwhile, we construct a novel additive similarity function following this paradigm, and expect to stimulate more valuable research. Further, we propose Convolutional Additive Token Mixer (CATM) employing underlying spatial and channel attention as novel interaction forms. This module eliminates troublesome complex operations such as matrix multiplication and Softmax. Finally, we evaluate our method on various vision tasks, and the throughput is reported on GPUs, ONNX and iPhone. Extensive experiments demonstrate our competitive performance compared with other SOTA backbones.
Related Work
Efficient Vision Transformers
@@
Since the debut of ViT (Dosovitskiy et al. 2020) and through its successful validation in image classification tasks on large-scale datasets such as ImageNet (Deng et al. 2009), it has demonstrated the potential of self-attention mechanisms for applications in computer vision (Dong et al. 2022). However, the concomitant increase in network sizes poses a huge challenge for resource-constrained scenarios such as mobile devices and real-time applications. To promote the potential value of ViT, workers have invested a considerable effort in efficient ViTs.
@@
The improvement ideas for ViT architecture cover several aspects. One of which is to refine the token mixer to enhance performance or solve the quadratic complexity problem of the self-attention mechanism. For example, PVT (Wang et al. 2021) adopts spatial reduction strategy to achieve sparse attention to deal with high-resolution images, and Swin (Liu et al. 2021) employs a window-splitting approach to achieve local self-attention and deals with dependencies between patches by window shifting.
@@
Another perspective is to explore hybrid models combining CNN and Transformer as compensation for the limitation of self-attention mechanism in processing local information (Guo et al. 2022). EdgeViT (Pan et al. 2022) adopts convolutional layers and sparse attention for information integration and propagation in block respectively. NextViT (Li et al. 2022a) confirms the validity of hybrid models and the design strategy through comprehensive experiments. EfficientViT (Liu et al. 2023) further analyzes the time-consuming proportion of various operations to achieve efficient inference.
Efficient Token Mixer
@@
Efficient token mixer design is one of the critical orientations in the evolution of Transformers. Efforts have been made to pursue lighter and more computationally efficient token mixers to advance the feasibility of training and practical applications.
@@
One part of the effort is devoted to improving the MSA. Twins (Chu et al. 2021) introduce self-attention between patches to achieve global and local dependency tandem. Works (Ali et al. 2021; Ding et al. 2022; Maaz et al. 2022) focus on information integration between channels and achieve self-attention between distinct domains through concatenation and transposition. Linear attention (Shen et al. 2021) addresses the quadratic complexity of the self-attention mechanism by assuming that the similarity function is linearly differentiable. (Han et al. 2023) promotes ReLU-based linear attention to be more discretized by focused function. (Bolya et al. 2022) further simplifies the attention complexity by increasing the attention heads.
@@
Heteromorphic-MSA (H-MSA) is an extension form of the self-attention mechanism development, which breaks through the limitations of MSA framework and intended to obtain better feature relations and more robust inference efficiency. Initially, MetaFormer (Yu et al. 2022) posited that token mixer is not the key component that affects the performance of Transformers, however PoolFormer could not be validated to be highly efficient. Subsequently, MobileViTv2 (Mehta and Rastegari 2022) simplifies the complex matrix multiplication by endowing global information to the context vector. SwiftFormer (Shaker et al. 2023) even eliminates ’Value’ and implements weighted sums between features with simpler normalization operations, achieving a more concise and efficient H-MSA.
Methods
@@
Overview of Self-Attention and Variants
@@
Separable self-attention (Sandler et al. 2018), see Fig. 2(b), reduces the matrix-based feature metric to a vector, achieving lightweight and efficient inference by decreasing the computational complexity. Subsequently, the context score is calculated through . Then, context score q is multiplied with K and summed along the spatial dimension to obtain the context vector that measures global information. It could be concretely described as:
@@
Swift self-attention, see Fig. 2(c), is a well-attended H-MSA architecture, which reduces the keys of self-attention to two, thereby achieving fast inference. It utilizes the coefficients obtained by linear transformed Q to weight each token. Subsequently, it is summed along the spatial domain and multiplied with K to obtain the global context. It is specifically represented as:
@@
Convolutional Additive Self-attention
@@
In this section, we argue that the information integration capability of self-attentive mechanisms resides in the multiple information interactions, e.g. the spatial association of MSA, DaViT (Ding et al. 2022) in channels, and the compressive representation of MobileViTv2 and SwiftFormer in both dimensions as in Fig. 2. Alternatively, isn’t it possible that simple and efficient operations can work better while satisfying multiple interactions?
Inheriting from this postulate, as shown in Fig. 2(d), we innovatively define the similarity function as the sum of context scores of and :
@@
where Query, Key, and Value are obtained by independent linear transformations, e.g. , , , denotes context mapping function, which contains the essential information interactions. The advantage of this generalization is that is not restricted to manual context design, and confers the possibility of implementing by convolutional operations. In this article, we simply concretize as the Sigmoid-based channel attention and spatial attention . Thus the output of CATM can be formulated as:
where denotes the linear transformation for integrating the contextual information. Since the operations in CATM are represented by convolution, the complexity is .
@@
Relation with Separable Self-Attention: Compared to (Mehta and Rastegari 2022)’s approach of extracting context scores only on Query, we perform similarity extraction on both Query and Key branches and retain the original feature dimensions on each branch. This enables greater preservation of visual sparse features and avoids information loss on 2D score vectors.
@@
Relation with Efficient Additive Attention: First, on token mixer, we adopt a sigmoid-activated form of attention extraction instead of normalization. this facilitates network parallelization and deployment on mobile devices. In addition, the attention module in (Shaker et al. 2023) is only applied to the last layer in each stage of the network, whereas the proposed CATM will be applied to each layer of the whole ViT architecture.
Complexity Analysis: In the concrete implementation, is designed as a combination of depthwise convolution and Sigmoid activation and , where denotes the batch size, which keeps constant in both training and inference phases. is achieved by a simplified channel attention and . Together with operations such as the mapping of QKV and the linear transformation of , CATM maintains linear complexity with respect to the input size:
Network Architecture
@@
Fig. 3(Upper) illustrates the proposed network architecture. Input a natural image of size . Subsequently it is downsampled to by two consecutive convolution layers with stride 2 in Stem.

### v2

@@ Abstract
Vision Transformers (ViTs) mark a revolutionary advance in neural networks with their token mixer’s powerful global context capability. However, the pairwise token affinity and complex matrix operations limit its deployment on resource-constrained scenarios and real-time applications, such as mobile devices, although considerable efforts have been made in previous works. In this paper, we introduce CAS-ViT: Convolutional Additive Self-attention Vision Transformers, to achieve a balance between efficiency and performance in mobile applications. Firstly, we argue that the capability of token mixers to obtain global contextual information hinges on multiple information interactions, such as spatial and channel domains. Subsequently, we propose Convolutional Additive Token Mixer (CATM) employing underlying spatial and channel attention as novel interaction forms. This module eliminates troublesome complex operations such as matrix multiplication and Softmax. We introduce Convolutional Additive Self-attention(CAS) block hybrid architecture and utilize CATM for each block. And further, we build a family of lightweight networks, which can be easily extended to various downstream tasks. Finally, we evaluate CAS-ViT across a variety of vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation. Our M and T model achieves 83.0%/84.1% top-1 with only 12M/21M parameters on ImageNet-1K. Meanwhile, throughput evaluations on GPUs, ONNX, and iPhones also demonstrate superior results compared to other state-of-the-art backbones. Extensive experiments demonstrate that our approach achieves a better balance of performance, efficient inference and easy-to-deploy. Our code and model are available at: https://github.com/Tianfang-Zhang/CAS-ViT
I Introduction
@@
In recent years, the emergence of Vision Transformers (ViTs) marks a revolutionary shift in neural network architecture [2, 3, 4]. Compared to Convolutional Neural Networks (CNNs), which make a splash with lower computational complexity and higher inference efficiency [5, 6], ViT provides advantages in capturing long-range dependencies, global modeling and representation. The basic module of ViT consists of token mixer [7, 8], MLP, and corresponding skip connections. In practice, token mixer is widely implemented as Multi-head Self-Attention (MSA), which operates on the entire input sequence, compensating for the limitations of CNNs with restricted receptive fields, and providing unique advantages in terms of model scalability and adaptability [9, 10, 11].
Nonetheless, deploying ViT models to mobile and edge devices remains a challenging task. Specifically, MSA is of quadratic complexity with respect to input image size [12, 13] leads to computational inefficiency that fails to meet the application requirements. Meanwhile, the complex matrix multiplication and Softmax operators in ViT suffer from operator unsupport and model conversion failure in deploying to edge devices [14]. As a result, on resource-constrained device deployment tasks or real-time applications, CNN is still the mainstream by virtue of its efficient computing and mature development support [15]. However, the restricted receptive field of CNN exposes problems such as insufficient feature extraction and underpowered performance. Therefore, developing token mixers that combine both efficient deployment and powerful performance in neural networks e.g. hybrid models, has become an urgent problem on mobile devices.
@@
Efforts have been made to improve token mixers, including MSA refinements and Heterogeneous-MSA (H-MSA). MSA refinements follow matrix multiplication of Query and Key. It emphasizes enhancing the ability to capture long-range dependencies under the self-attention mechanism while reducing the algorithm complexity. Specific techniques include feature shift [13], carrier token [16], sparse attention [17], linear attention [18], etc. Furthermore, H-MSA is an evolutionary variant that breaks the restriction of matrix multiplication of Q and K to explore more flexible network designs [7]. The recently proposed pool token mixer [7] and context vector [19, 20] further accelerate the inference efficiency.
@@
Although considerable progress has been made in recent works, they still suffer from the following constraints: 1. High complexity associated with matrix operators such as Softmax and multiplication in token mixers limits efficient inference, while remaining problematic for mobile deployments. 2. Networks composed of low-complexity operators fail to meet the increasing demand for powerful performance on real-time devices. The efforts still fail to achieve a satisfactory balance between efficiency and accuracy.
In this paper, we propose a family of lightweight networks named as Convolutional Additive Self-attention (CAS)-ViT to address the encountered problems. In particular, we argue that the capability of token mixers to obtain global contextual information hinges on multiple information interactions, such as spatial and channel domains. Further, by constructing an easy-to-implement additive similarity function, we eliminate computationally expensive operations and achieve self-attentive convolutional substitution. The main contributions of our work are summarized as follows:
1.
We propose Convolutional Additive Token Mixer (CATM) employing underlying spatial and channel attention as novel interaction forms. This module eliminates troublesome complex operations such as matrix multiplication and Softmax.
2.
We propose Convolutional Additive Self-attention(CAS) block hybrid architecture and utilize an H-MSA structure known as CATM for each block. Further we build a family of lightweight networks called CAS-ViT, which is a powerful backbone and can be extended to various downstream tasks.
3.
We evaluate our method on various vision tasks, subsequently deploy it on mobile devices and report throughput data. Our XXS model achieves 78.7% top-1 on ImageNet-1K and 14.6 throughput on iPhone. Moreover, the Tiny model achieves 84.1% top-1 with only 21M parameters, as shown in Fig.1. Our models outperform recent popular works like SwiftFormer [20], MobileViTv2 [21]. Extensive experiments demonstrate that our approach achieves a better balance of performance, efficient inference and easy-to-deploy.
II Related Work
II-A Efficient Vision Transformers
@@
In recent years, ViTs have represented a pivotal shift in computer vision, and its research has shown a remarkable evolution [22, 23]. Since the debut of ViT [24] and through its successful validation in image classification tasks on large-scale datasets such as ImageNet [1], it has demonstrated the potential of self-attention mechanisms for applications in computer vision [25, 26]. However, the concomitant increase in network sizes poses a huge challenge for resource-constrained scenarios such as mobile devices and real-time applications. To promote the potential value of ViT, workers have invested a considerable effort in efficient ViTs [27, 28, 29, 30].
@@
The improvement ideas for ViT architecture cover several aspects. One of which is to refine the token mixer to enhance performance or solve the quadratic complexity problem of the self-attention mechanism. For example, PVT [31] adopts spatial reduction strategy to achieve sparse attention to deal with high-resolution images, and Swin [13] employs a window-splitting approach to achieve local self-attention and deals with dependencies between patches by window shifting.
@@
Another perspective is to explore hybrid models combining CNN and Transformer as compensation for the limitation of the self-attention mechanism in processing local information [32]. EdgeViT [33] adopts convolutional layers and sparse attention for information integration and propagation in block respectively. NextViT [34] confirms the validity of hybrid models and the design strategy through comprehensive experiments. EfficientViT [29] further analyzes the time-consuming proportion of various operations to achieve efficient inference. Efficient ViTs have demonstrated exceptional versatility, excelling not only in image classification but also in a wide range of visual tasks, laying the foundation for future development.
II-B Efficient Token Mixer
@@
Efficient token mixer design is one of the critical orientations in the evolution of Transformers [7, 8]. Efforts have been made to pursue lighter and more computationally efficient token mixers to advance the feasibility of training and practical applications [35, 36, 37].
@@
One part of the effort is devoted to refining MSA by following the principles of Q and K matrix multiplication. Twins [38] introduce self-attention between patches to achieve global and local dependency tandem. Works [39, 40, 41] focus on information integration between channels and achieve self-attention between distinct domains through concatenation and transposition. Linear attention [28] addresses the quadratic complexity of the self-attention mechanism by assuming that the similarity function is linearly differentiable. [18] promotes ReLU-based linear attention to be more discretized by focused function. [42] further simplifies the attention complexity by increasing the attention heads.
@@
Heteromorphic-MSA (H-MSA) is an extension form of the self-attention mechanism development, which streamlines the multiply operations in MSA framework and intended to obtain robust features and more efficient inference performance. Initially, MetaFormer [7] posited that token mixer is not the key component that affects the performance of Transformers, however PoolFormer could not be validated to be highly efficient. Subsequently, MobileViTv2 [21] simplifies the complex matrix multiplication by endowing global information to the context vector. SwiftFormer [20] even eliminates ’Value’ and implements weighted sums between features with simpler normalization operations, achieving a more concise and efficient H-MSA. Inspired by previous excellent work, we rethink the token mixer design and argue that the facilitation requires information interaction on spatial and channel domains. Detailed work is described as follows.
III Methods
@@
III-A Overview of Self-Attention and Variants
@@
Separable self-attention [19], see Fig. 2(b), reduces the matrix-based feature metric to a vector, achieving lightweight and efficient inference by decreasing the computational complexity. Subsequently, the context score is calculated through . Then, context score q is multiplied with K and summed along the spatial dimension to obtain the context vector that measures global information. It could be concretely described as:
@@
Swift self-attention [20], see Fig. 2(c), is a well-attended H-MSA architecture, which reduces the keys of self-attention to two, thereby achieving fast inference. It utilizes the coefficients obtained by linear transformed Q to weight each token. Subsequently, it is summed along the spatial domain and multiplied with K to obtain the global context. It is specifically represented as:
@@
III-B Convolutional Additive Token Mixer
Motivation: The attention graph is obtained by measuring all Query-pairs in MSA as Fig. 2(a), which leads to quadratic computational complexity . H-MSA approaches expect to measure the correlation between tokens through a context vector as Fig. 2(b)(c). The concept of context vector is inherited from MSA and simplified, and its generality is that it requires to perform interaction between spatial and channel domains, which can be confirmed by the operation of and vectors in Fig. 2.
@@ III-B Convolutional Additive Token Mixer
Motivated by this, we argue that the information integration capability of self-attentive mechanisms resides in the multiple information interactions, e.g. the spatial association of MSA, DaViT [40] and Twins [38] in channels, and the compressive context vector of MobileViTv2 [21] and SwiftFormer [20] in both dimensions as in Fig. 2. Alternatively, isn’t it possible that simple and efficient operations can work better while satisfying multiple interactions?
Token Mixer Design: Inheriting from this postulate, we can consider simpler operators to meet the requirements, rather than getting involved in matrix multiplication that imposes an inference burden or Softmax that imposes a risk of not being deployable. Along this thought, for a tensor of size , the multiple information interactions can be set to spatial () and channel () domains. Indeed, other studies on scale decomposition have used multiple heads on channels to parallelize computation as well as dividing patches on spatial domain for speedup, however, these techniques are subdivided on the base dimensions for feature tensor.
We consider a spatial interaction as in Fig. 3 (Right) with input . The local token information is first integrated by a 33 convolution layer, after which the dimension is reduced to 1 with a 11 convolution layer, followed by a Sigmoid activation as an attention map in the spatial domain. Specifically represented as:
For channel domain interaction, we refer to SENet [43], but instead of channel reduction we use 11 convolution to integrate the information between channels:
where denotes the adaptive pooling and is the group convolution layer, group number is set to channel number by default. Stacking these two operations yields the feature map after spatial and channel domain interactions, denoted as .
As shown in Fig. 2(d), we innovatively define the similarity function as the sum of context scores of :
@@
where Query, Key, and Value are obtained by independent linear transformations, e.g. , , , denotes context mapping function, which contains the essential information interactions. The advantage of this generalization is that it is not restricted to manual context design, and confers the possibility of implementing by convolutional operations. Performing multiple interactions on both Q and K is expected that the network learns and notices more valuable tokens. The two are integrated through additive operations, a linear approach that avoids the high complexity of matrix multiplication and is also effective in retaining useful information. Thus the output of CATM can be formulated as:
where denotes the linear transformation for integrating the contextual information.
Analysis of Receptive Field: ViT [24] would perform MSA after dividing the original image into patches to avoid high computational load, which would limit the receptive field to inside the patches e.g. 77. There are many works devoted to generalize the receptive field to global [13]. In fact, a 77 receptive field inside the patch can be realized by a simple three-layer 33 convolutional stacking, although it is different from the correlation and receptive field of the self-attention mechanism, a similar effect can be achieved by multiple information interactions and convolutional stacking.
@@
Relation with Separable Self-Attention: Compared to [21]’s approach of extracting context scores only on Query, we perform similarity extraction on both Query and Key branches and retain the original feature dimensions on each branch. This enables greater preservation of visual sparse features and avoids information loss on 2D score vectors.
@@
Relation with Efficient Additive Attention: First, on token mixer, we adopt a sigmoid-activated form of attention extraction instead of normalization. this facilitates network parallelization and deployment on mobile devices. In addition, the attention module in [20] is only applied to the last layer in each stage of the network, whereas the proposed CATM will be applied to each layer of the whole ViT architecture.
Complexity Analysis: In the concrete implementation, spatial and channel domain interactions are designed as a combination of depthwise convolution and Sigmoid activation, where the convolution kernel size was set to 33, i.e:
The operation in CATM consists of four parts: the QKV separation convolution, and and the final mapping layer . CATM maintains linear complexity with respect to the input size:
III-C Convolutional Additive Self-attention Block
CAS block serves as the basic structure of our method and is stacked in each stage. The design of CAS block is referenced to the widely validated hybrid model[29, 33] and contains three parts with residual shortcuts: Integration subnet, CATM and MLP, as shown in Fig. 3 (Left Lower).
The integration sub-network consists of three 33 depth-wise convolution layers activated by ReLU [44], this module increases the receptive field, integrates local information to facilitate the subsequent CATM, and acts as the positional encoding layer. CATM is utilized in all blocks at each stage of the network, which is a simplified design that facilitates future development. Since the underlying network is implemented as a convolution, the CAS block can strike a good balance between computational efficiency and deployment. MLP ratio was set to 4 by default.
III-D Network Architecture
@@
Fig. 3(Left Upper) illustrates the proposed network architecture. Input a natural image of size . It is downsampled to by two consecutive 33 convolution layers with stride 2 in Stem.